Question 1 gives you the information of the stocks and their correlation, which will be used in many of the following questions.

James, a portfolio manager, would like to form the following portfolio between Microsoft and Coca-Cola:

            Expected return (%)      Standard Deviation (%) Weight

Microsoft                 28                          42                    0.4                  

Coca-Cola              12.5                         21                    0.6

The correlation between the two stocks is -0.2.

What is the standard deviation of the portfolio? ______%
Answer #1

wiai+wia+2w, W2 Pi 2 C1 G2 oportfolio Where: Proportion of the portfolio invested in Asset 1 Proportion of the portfolio inve

\sigma = \sqrt{(0.4 * 0.28)^2 + (0.6 * 0.125)^2 + (2 * 0.4 * 0.6 * 0.28 * 0.125 * -0.2)}

\sigma = \sqrt{0.028224 + 0.015876 - 0.00847}

\sigma = \sqrt{0.035633}

\sigma = 0.1888

Standard deviation = 18.88%